OpenStudy (anonymous):

Solve the equation. |-13| + 8(2) = 30y - |-35| - 29y ^ the 2 is one top right corner of the 8

OpenStudy (mertsj):

\[|-13|+8^2=30y-|-35|-29y\]

OpenStudy (mertsj):

\[13+64=30y-35-29y\]

OpenStudy (mertsj):

\[77=y-35\]

OpenStudy (anonymous):

are you sure beacuse its either A.42 B.-112 C.112 D.-42

OpenStudy (mertsj):

\[112=y\]